# Generalize across Homophily and Heterophily: Hybrid Spectral Graph Pre-Training and Prompt Tuning

### v2

@@ B.2 Implementation Details
For the baseline implementations, including GCN, GAT, DGI, GraphCL, SimGRACE, GPPT, GPrompt, GPF-plus, and All-in-One, we leveraged the public benchmark ProG (Zi et al. 2024) to implement them using both 2-layer GCN and UniMP (a Graph Transformer model) as the backbone GNNs. For BernNet, ChebNetII, TFE-GNN, PolyGCL, and ProNoG, we use the source code published by the original authors. For baselines involving a pre-training stage, we adopted the hyperparameters recommended in the respective source code to pre-train the models. During the downstream tuning phase, hyperparameters such as the learning rate and epochs were kept consistent with those of our proposed methods to ensure a fair comparison.
For homophilic graph prompt baselines implemented via ProG, we follow the ProG benchmark to perform a grid search for optimal pretraining methods on each dataset. Specifically, we evaluate six pretraining methods: DGI (Veličković et al. 2018), GraphMAE (Hou et al. 2022), EdgePreGPPT (Sun et al. 2022), EdgePreGprompt (Liu et al. 2023), GraphCL (You et al. 2020), and SimGRACE (Xia et al. 2022). For the heterophilic graph prompt baseline ProNoG, we adopt the two pretraining methods recommended in its original paper: DGI (Veličković et al. 2018) and GraphCL (You et al. 2020). We report the best results from combinations of pretraining methods and prompts for each dataset. Optimal pretraining-prompt tuning combinations are detailed in Tables 9, 10, and 11.
@@ B.3 Implementation Details
  • Cora (McCallum et al. 2000), Citeseer (Sen et al. 2008), and Pubmed (Kipf and Welling 2016) are three citation networks that are regarded as typical homophilic graphs. Within these networks, the nodes stand for papers, and the edges signify the citation connections between two papers. The features are comprised of bag-of-word representations of the papers, and the labels serve to denote the research topic of each paper.
  • Cornell, Texas, and Wisconsin (Pei et al. 2020) are three heterophilic networks that come from the WebKB project. In these networks, the nodes are web pages of computer science departments of different universities, and the edges are hyperlinks between them. The features of each page are represented as bag-of-words, and the labels indicate the types of web pages.
  • Chameleon and Squirrel (Rozemberczki, Allen, and Sarkar 2021) are two heterophilic networks built upon Wikipedia. In these networks, the nodes represent web pages within Wikipedia, and the edges represent the links that connect them. The features are composed of informative nouns found on the Wikipedia pages, and the labels show the average traffic that the web pages receive.
  • Amazon-ratings and Roman-empire (Platonov et al. 2023) are substantial heterophilic graphs that possess distinct structural properties and originate from different fields. These graphs are proposed to mitigate the issues present in existing heterophilic graphs. For instance, the networks of Chameleon and Squirrel contain a significant number of duplicate nodes, which results in problems such as training and test data leakage.
@@ B.4 Synthetic Dataset Details
We utilize the widely-employed Contextual Stochastic Block Model (CSBM) (Chien et al. 2020) to generate synthetic datasets for validating our theorems. Specifically, these graphs are distinguished by adjustable edge probabilities both within and between different classes. The fundamental concept is that nodes within the same class display a uniform feature distribution. The graph is generated as $\mathcal{G}\sim CSBM(n,f,\sigma,\mu)$, where $n$ stands for the number of nodes, $f$ represents the feature dimension and $\sigma$ and $\mu$ are hyperparameters that affect the graph structure and node features, respectively.
@@
Here $d$ is the average degree of nodes. Following previous research (Wan et al. 2024), we can adjust the homophily level $h$ by modifying $\sigma=\sqrt{d(2h-1)}$, thereby obtaining graphs with different structures. For our validation experiments, we set $d=50$, $n=3000$, and $f=128$, and generate synthetic graphs with the homophily level $h$ ranging from 0 to 1.
@@ B.5 Experiments on Synthetic Datasets
To present our theoretical findings in a more intuitive way, we further conduct the validation experiments to verify Theorem 3.3. We utilize GNNs with different filtering characteristics to learn the CSBM graphs graph generated in Figure 1. To avoid complex calculations, following previous work (Luo et al. 2024), we use functions with linear or quadratic terms to construct the graph filter. We adopt three graph filters: low-pass, mid-pass, and high-pass filters, which can be formulated as:
@@
We apply local-global contrastive learning (Hassani and Khasahmadi 2020) and subsequently evaluate the performance through a node classification task. To present the performance differences of different graph filters more clearly, we set the ratio of the training set, validation set, and test set to 5:2:3. The visualization and results of the three filters are shown in Figure 5 and are consistent with Theorem 3.3, indicating that a single graph filter cannot generalize across graphs with varying spectral distributions.
@@ C.2 Algorithm of HS-GPPT
Output: The pre-trained graph fitlers $g_{k,C-k}$ and its corresponding integration weights $\mathbf{w}^{k}$.
@@ C.3 Complexity Analysis
For a $C^{\prime}$-layer GCN (Kipf and Welling 2016), the parameter complexity is $O(C^{\prime}d^{2})$. In our pre-training stage, the tuned parameter is $O(C^{\prime}d^{2}+C^{\prime}d)$. Due to the fact that $C^{\prime}\ll d$, our pre-trained parameters are lightweight. In the prompt-tuning stage, we freeze the pre-trained graph filters and integration weights, so the tuned parameter is $O(C^{\prime}N_{p}d)$. Since both $C^{\prime}\ll d$ and $N_{p}\ll d$, our prompt tuning is also lightweight.
@@ D.1 Graph prompt tuning
Graph Neural Networks (GNNs) (Kipf and Welling 2016; Defferrard, Bresson, and Vandergheynst 2016; Veličković et al. 2017) have emerged as a powerful framework for learning from graph data, which enable information propagation and feature extraction through iterative message passing among neighboring nodes. Typically, GNNs operate under a supervised setting, where models are trained for a specific task on the input graph and make inferences on the same graph. However, the difficulty of obtaining labeled data (Chen et al. 2024) limits the performance. To break through this limitation, some research (Jin et al. 2020; You et al. 2020; Xia et al. 2022; Zhu et al. 2021; Hou et al. 2023; Xia et al. 2023) have turned to self-supervised learning on graph data. These methods first pre-train a graph model using self-supervised tasks, and then fine-tune it with downstream tasks. This approach enables the pre-trained knowledge to enhance the performance of the downstream tasks. Nevertheless, it is proved that the misalignment between the pretexts and downstream tasks will impede knowledge transfer and even lead to negative transfer (Wang et al. 2021).
In response to this challenge, graph prompt tuning (Sun et al. 2022; Fang et al. 2024; Sun et al. 2023; Liu et al. 2023; Yu et al. 2024) has emerged as an attractive alternative to the conventional fine-tuning paradigm. The core idea is to design prompts to manipulate the downstream tasks, allowing them to better align with the frozen pre-trained model. For example, GPrompt (Liu et al. 2023) uses prompt vectors to unify pre-training and downstream tasks under a common template, while GPF (Fang et al. 2024) inserts prompt nodes adaptable to various pre-training strategies. However, existing methods are primarily designed for homophilic graphs and largely limited to low-frequency information. This is due to their reliance on low-pass graph filters (e.g., GCN (Kipf and Welling 2016)) and homophilic pre-training tasks such as link prediction (Sun et al. 2022; Liu et al. 2023), which are known to retain mainly low-frequency signals (Liu et al. 2022; Yu et al. 2024). As a result, they can be viewed as special cases of our approach when restricted to low-frequency components.
Our spectral analysis reveals that in settings with sparse supervision, large spectral gaps hinder effective alignment and parameter learning, which contributes to the failure of these homophily-based methods. Although ProNoG (Yu et al. 2024) takes a first step toward heterophilic prompt tuning, it lacks a comprehensive spectral perspective. In contrast, our method enriches pre-trained knowledge, reduces spectral gaps, and enables effective alignment and robust performance even with limited labels.
@@ D.2 Heterophilic Graph Learning
To address the heterophilic issue, many methods have been proposed, such as capturing high-frequency information (Bo et al. 2021), discovering potential neighbors (Jin et al. 2021; Pei et al. 2020), and engaging in high-order message passing (Zhu et al. 2020). While they are trained under the supervised setting, there are also methods (He et al. 2023; Xiao et al. 2022, 2024; Chen, Lei, and Wei 2024; Wan et al. 2024) further aiming to design self-supervised learning for heterophilic graphs. These methods attempt to adopt more diverse neighbor selection approaches or learn high-frequency information (Chen, Lei, and Wei 2024; Wan et al. 2024) through self-supervised tasks. However, these methods are tailored for full-model fine-tuning and do not provide a mechanism to transfer this knowledge to downstream tasks through lightweight prompts. Therefore, directly using the heterophilic priors can cause negative transfer, especially under limited supervision. Although graph prompt tuning holds the promise of enabling the transfer of pre-trained knowledge, the application of this paradigm to heterophilic graphs has been scarcely investigated. Our research commences in the spectral domain and conducts an in-depth understanding and exploration of graph prompt tuning specifically for heterophilic graphs.
Appendix E Experiment Results and Limitations
In this section, we first elaborate on the limitations of our research, followed by the presentation of the complete experimental results.
E.1 Limitations
Our study achieves promising results on commonly used graph datasets (e.g., Cora, Citeseer, Pubmed), which are valuable for benchmarking but may not fully reflect the variability of real-world graph environments, such as those with sparse edges, dynamic interactions, or domain-specific features in areas like urban mobility or social media. While our approach provides a robust foundation for spectral alignment and prompt tuning, exploring its performance on more diverse real-world datasets—particularly those with complex noise or evolving structures—could further strengthen its practical relevance. Future efforts might also involve adapting the framework to handle specialized scenarios, such as integrating domain knowledge or improving efficiency for large-scale graphs, to broaden its applicability across different contexts.
E.2 Parameter Sensitivity
@@
(i) Regarding the threshold for cross edges, we set the inner edge threshold at 0.2 (in line with the previous work (Sun et al. 2023)) and vary the cross edge threshold from 0 to 1. The model is evaluated on three datasets with varying degrees of homophily: Cora, Wisconsin, and Empire. The results reveal a distinct homophily-driven heuristic in the selection of the optimal threshold: the optimal cross edge threshold for homophilic graphs (e.g., Pubmed) is higher than that for heterophilic graphs (e.g., Wisconsin and Empire). This systematic variation stems from our dual objectives of maintaining structural integrity and ensuring the effectiveness of the prompts. In homophilic graphs, where neighboring nodes tend to be similar, we opt for a higher $\tau_{cross}$ value to ensure that connected nodes have more similar features and preserve the graph’s homophilic nature. Conversely, in heterophilic graphs with more dissimilar neighboring nodes, we set a lower $\tau_{cross}$ to facilitate more diverse connections between prompt and original graph nodes.
@@ E.3 Runtime Analysis
E.4 Additional Accuracy Results
@@
GCN
0.6532
$\pm$0.0319
0.6506
$\pm$0.0185
0.5546
$\pm$0.0367
0.5405
$\pm$0.0358
0.4442
$\pm$0.0303
0.4187
$\pm$0.0220
0.2384
$\pm$0.0150
0.1911
$\pm$0.0634
0.1613
$\pm$0.0145
0.1491
$\pm$0.0707
GAT
0.6018
$\pm$0.0238
0.6091
$\pm$0.0201
0.5472
$\pm$0.0641
0.5381
$\pm$0.0595
0.4446
$\pm$0.0350
0.4162
$\pm$0.0332
0.2002
$\pm$0.0248
0.1827
$\pm$0.0371
0.1419
$\pm$0.0221
0.1367
$\pm$0.0208
BernNet
0.6494
$\pm$0.0280
0.6356
$\pm$0.0296
0.6584
$\pm$0.0216
0.6493
$\pm$0.0215
0.4888
$\pm$0.0197
0.4712
$\pm$0.0192
0.2581
$\pm$0.0074
0.2030 $\pm$0.0063
0.1215
$\pm$0.0280
0.1071
$\pm$0.0253
ChebNetII
0.6798
$\pm$0.0170
0.6752
$\pm$0.0202
0.6616
$\pm$0.0431
0.6554
$\pm$0.0410
0.4526
$\pm$0.0353
0.4304
$\pm$0.0328
0.2382
$\pm$0.0148
0.1970
$\pm$0.0047
0.0661
$\pm$0.0088
0.0551
$\pm$0.0054
TFE-GNN
0.3376
$\pm$0.0198
0.3286
$\pm$0.0177
0.4694
$\pm$0.0170
0.4590
$\pm$0.0155
0.2932
$\pm$0.0237
0.2874
$\pm$0.0203
0.2301
$\pm$0.0099
0.1939
$\pm$0.0014
0.3502
$\pm$0.0151
0.3155
$\pm$0.0182
DGI
0.3658
$\pm$0.1487
0.3187
$\pm$0.1251
0.4870
$\pm$0.0834
0.3743
$\pm$0.1073
0.2922
$\pm$0.0347
0.2235
$\pm$0.0405
0.1818
$\pm$0.0418
0.1485
$\pm$0.0237
0.1209
$\pm$0.0359
0.0710
$\pm$0.0188
GraphCL
0.5958
$\pm$0.0395
0.5603
$\pm$0.0486
0.5694
$\pm$0.0327
0.5576
$\pm$0.0396
0.4154
$\pm$0.0263
0.3741
$\pm$0.0356
0.2326
$\pm$0.0358
0.1853
$\pm$0.0158
0.1184
$\pm$0.0245
0.0961
$\pm$0.0150
SimGRACE
0.4400
$\pm$0.0321
0.4283
$\pm$0.0341
0.4416
$\pm$0.0424
0.4316
$\pm$0.0482
0.3664
$\pm$0.0392
0.3412
$\pm$0.0405
0.2064
$\pm$0.0239
0.1761
$\pm$0.0153
0.0878
$\pm$0.0097
0.0635
$\pm$0.0065
PolyGCL
0.6948
$\pm$0.0284
0.6655
$\pm$0.0221
0.6792
$\pm$0.0364
0.6782
$\pm$0.0357
0.5266 $\pm$0.0286
0.5010
$\pm$0.0224
0.2048
$\pm$0.0155
0.1874
$\pm$0.0081
0.0634
$\pm$0.0060
0.0528
$\pm$0.0029
GPPT
0.7026
$\pm$0.0172
0.6611
$\pm$0.0313
0.6540
$\pm$0.0452
0.6459
$\pm$0.0447
0.3854
$\pm$0.0617
0.3731
$\pm$0.0608
0.2345
$\pm$0.0156
0.1944
$\pm$0.0128
0.0750
$\pm$0.0037
0.0658
$\pm$0.0036
GPrompt
0.6578
$\pm$0.0323
0.6366
$\pm$0.0252
0.6230
$\pm$0.0428
0.6113
$\pm$0.0428
0.4568
$\pm$0.0269
0.4316
$\pm$0.0315
0.2269
$\pm$0.0252
0.1946
$\pm$0.0125
0.0652
$\pm$0.0071
0.0539
$\pm$0.0040
GPF-plus
0.6250
$\pm$0.0279
0.6160
$\pm$0.0251
0.6706
$\pm$0.0242
0.6534
$\pm$0.0250
0.4564
$\pm$0.0308
0.4441
$\pm$0.0322
0.2092
$\pm$0.0273
0.1826
$\pm$0.0177
0.0655
$\pm$0.0080
0.0476
$\pm$0.0055
All-in-One
0.5442
$\pm$0.0731
0.4308
$\pm$0.0410
0.6724
$\pm$0.0335
0.6600
$\pm$0.0330
0.4616
$\pm$0.0095
0.3638
$\pm$0.0158
0.2679 $\pm$0.0371
0.1819
$\pm$0.0263
0.0566
$\pm$0.0037
0.0472
$\pm$0.0028
ProNoG
0.5838
$\pm$0.0121
0.5564
$\pm$0.0203
0.5348
$\pm$0.0748
0.5242
$\pm$0.0855
0.2596
$\pm$0.0160
0.2466
$\pm$0.0172
0.2513
$\pm$0.0426
0.1963
$\pm$0.0102
0.1025
$\pm$0.0142
0.0784
$\pm$0.0090
HS-GPPT 0.7156 $\pm$0.0132 0.6915 $\pm$0.0166 0.6912 $\pm$0.0215 0.6910 $\pm$0.0199
0.5188
$\pm$0.0351
0.5043 $\pm$0.0223
0.2407
$\pm$0.0239
0.1972
$\pm$0.0123
0.3836 $\pm$0.0322 0.3520 $\pm$0.0194

### v3

@@ B.2 Implementation Details
For the baseline implementations, including GCN, GAT, DGI, GraphCL, SimGRACE, GPPT, GPrompt, GPF-plus, and All-in-One, we leveraged the public benchmark ProG Zi et al. [2024] to implement them using both 2-layer GCN and UniMP (a Graph Transformer model) as the backbone GNNs. For BernNet, ChebNetII, TFE-GNN, PolyGCL, and ProNoG, we use the source code published by the original authors. For baselines involving a pre-training stage, we adopted the hyperparameters recommended in the respective source code to pre-train the models. During the downstream tuning phase, hyperparameters such as the learning rate and epochs were kept consistent with those of our proposed methods to ensure a fair comparison.
For homophilic graph prompt baselines implemented via ProG, we follow the ProG benchmark to perform a grid search for optimal pretraining methods on each dataset. Specifically, we evaluate six pretraining methods: DGI Veličković et al. [2018], GraphMAE Hou et al. [2022], EdgePreGPPT Sun et al. [2022], EdgePreGprompt Liu et al. [2023], GraphCL You et al. [2020], and SimGRACE Xia et al. [2022]. For the heterophilic graph prompt baseline ProNoG, we adopt the two pretraining methods recommended in its original paper: DGI Veličković et al. [2018] and GraphCL You et al. [2020]. We report the best results from combinations of pretraining methods and prompts for each dataset. Optimal pretraining-prompt tuning combinations are detailed in Tables 9, 10, and 11.
@@ B.3 Implementation Details
  • Cora McCallum et al. [2000], Citeseer Sen et al. [2008], and Pubmed Kipf and Welling [2016] are three citation networks that are regarded as typical homophilic graphs. Within these networks, the nodes stand for papers, and the edges signify the citation connections between two papers. The features are comprised of bag-of-word representations of the papers, and the labels serve to denote the research topic of each paper.
  • Cornell, Texas, and Wisconsin Pei et al. [2020] are three heterophilic networks that come from the WebKB project. In these networks, the nodes are web pages of computer science departments of different universities, and the edges are hyperlinks between them. The features of each page are represented as bag-of-words, and the labels indicate the types of web pages.
  • Chameleon and Squirrel Rozemberczki et al. [2021] are two heterophilic networks built upon Wikipedia. In these networks, the nodes represent web pages within Wikipedia, and the edges represent the links that connect them. The features are composed of informative nouns found on the Wikipedia pages, and the labels show the average traffic that the web pages receive.
  • Amazon-ratings and Roman-empire Platonov et al. [2023] are substantial heterophilic graphs that possess distinct structural properties and originate from different fields. These graphs are proposed to mitigate the issues present in existing heterophilic graphs. For instance, the networks of Chameleon and Squirrel contain a significant number of duplicate nodes, which results in problems such as training and test data leakage.
@@ B.4 Synthetic Dataset Details
We utilize the widely-employed Contextual Stochastic Block Model (CSBM) Chien et al. [2020] to generate synthetic datasets for validating our theorems. Specifically, these graphs are distinguished by adjustable edge probabilities both within and between different classes. The fundamental concept is that nodes within the same class display a uniform feature distribution. The graph is generated as $\mathcal{G}\sim CSBM(n,f,\sigma,\mu)$, where $n$ stands for the number of nodes, $f$ represents the feature dimension and $\sigma$ and $\mu$ are hyperparameters that affect the graph structure and node features, respectively.
@@
Here $d$ is the average degree of nodes. Following previous research Wan et al. [2024], we can adjust the homophily level $h$ by modifying $\sigma=\sqrt{d(2h-1)}$, thereby obtaining graphs with different structures. For our validation experiments, we set $d=50$, $n=3000$, and $f=128$, and generate synthetic graphs with the homophily level $h$ ranging from 0 to 1.
@@ B.5 Experiments on Synthetic Datasets
To present our theoretical findings in a more intuitive way, we further conduct the validation experiments to verify Theorem 3.3. We utilize GNNs with different filtering characteristics to learn the CSBM graphs graph generated in Figure 1. To avoid complex calculations, following previous work Luo et al. [2024], we use functions with linear or quadratic terms to construct the graph filter. We adopt three graph filters: low-pass, mid-pass, and high-pass filters, which can be formulated as:
@@
We apply local-global contrastive learning Hassani and Khasahmadi [2020] and subsequently evaluate the performance through a node classification task. To present the performance differences of different graph filters more clearly, we set the ratio of the training set, validation set, and test set to 5:2:3. The visualization and results of the three filters are shown in Figure 5 and are consistent with Theorem 3.3, indicating that a single graph filter cannot generalize across graphs with varying spectral distributions.
@@ C.2 Algorithm of HS-GPPT
Output: The pre-trained graph filters $g_{k,C-k}$ and its corresponding integration weights $\mathbf{w}^{k}$.
@@ C.3 Complexity Analysis
For a $C^{\prime}$-layer GCN Kipf and Welling [2016], the parameter complexity is $O(C^{\prime}d^{2})$. In our pre-training stage, the tuned parameter is $O(C^{\prime}d^{2}+C^{\prime}d)$. Due to the fact that $C^{\prime}\ll d$, our pre-trained parameters are lightweight. In the prompt-tuning stage, we freeze the pre-trained graph filters and integration weights, so the tuned parameter is $O(C^{\prime}N_{p}d)$. Since both $C^{\prime}\ll d$ and $N_{p}\ll d$, our prompt tuning is also lightweight.
@@ D.1 Graph prompt tuning
Graph Neural Networks (GNNs) Kipf and Welling [2016]; Defferrard et al. [2016]; Veličković et al. [2017] have emerged as a powerful framework for learning from graph data, which enable information propagation and feature extraction through iterative message passing among neighboring nodes. Typically, GNNs operate under a supervised setting, where models are trained for a specific task on the input graph and make inferences on the same graph. However, the difficulty of obtaining labeled data Chen et al. [2024b] limits the performance. To break through this limitation, some research Jin et al. [2020]; You et al. [2020]; Xia et al. [2022]; Zhu et al. [2021]; Hou et al. [2023]; Xia et al. [2023] have turned to self-supervised learning on graph data. These methods first pre-train a graph model using self-supervised tasks, and then fine-tune it with downstream tasks. This approach enables the pre-trained knowledge to enhance the performance of the downstream tasks. Nevertheless, it is proven that the misalignment between the pretexts and downstream tasks will impede knowledge transfer and even lead to negative transfer Wang et al. [2021].
In response to this challenge, graph prompt tuning Sun et al. [2022]; Fang et al. [2024]; Sun et al. [2023]; Liu et al. [2023]; Yu et al. [2024] has emerged as an attractive alternative to the conventional fine-tuning paradigm. The core idea is to design prompts to manipulate the downstream tasks, allowing them to better align with the frozen pre-trained model. For example, GPrompt Liu et al. [2023] uses prompt vectors to unify pre-training and downstream tasks under a common template, while GPF Fang et al. [2024] inserts prompt nodes adaptable to various pre-training strategies. However, existing methods are primarily designed for homophilic graphs and are largely limited to low-frequency information. This is due to their reliance on low-pass graph filters (e.g., GCN Kipf and Welling [2016]) and homophilic pre-training tasks such as link prediction Sun et al. [2022]; Liu et al. [2023], which are known to retain mainly low-frequency signals Liu et al. [2022]; Yu et al. [2024]. As a result, they can be viewed as special cases of our approach when restricted to low-frequency components.
Our spectral analysis reveals that in settings with sparse supervision, large spectral mismatch hinders effective alignment and parameter learning, which contributes to the failure of these homophily-based methods. While ProNoG Yu et al. [2024], HeterGP Yan et al. [2025], and DAGPrompT Chen et al. [2025] have recently extended prompt tuning to heterophilic graphs via multi-view or distribution-aware strategies, they primarily operate in the spatial domain or rely on heuristic adaptations. Crucially, they overlook the fundamental spectral mismatch between the frozen backbone and diverse downstream graphs. In contrast, our method explicitly treats prompt tuning as a spectral alignment process, enriching pre-trained knowledge and actively aligning the input spectrum to match pre-trained filters, ensuring robust performance even with limited labels.
@@ D.2 Heterophilic Graph Learning
To address the heterophilic issue, many methods have been proposed, such as capturing high-frequency information Bo et al. [2021], discovering potential neighbors Jin et al. [2021]; Pei et al. [2020], and engaging in high-order message passing Zhu et al. [2020]. While they are trained under the supervised setting, there are also methods Xiao et al. [2024]; Chen et al. [2024a]; Wan et al. [2024] further aiming to design self-supervised learning for heterophilic graphs. These methods attempt to adopt more diverse neighbor selection approaches or learn high-frequency information Chen et al. [2024a]; Wan et al. [2024] through self-supervised tasks. However, these methods are tailored for full-model fine-tuning and do not provide a mechanism to transfer this knowledge to downstream tasks through lightweight prompts. Therefore, directly using the heterophilic priors can cause negative transfer, especially under limited supervision. Although graph prompt tuning holds the promise of enabling the transfer of pre-trained knowledge, the application of this paradigm to heterophilic graphs has been scarcely investigated. Our research commences in the spectral domain and conducts an in-depth understanding and exploration of graph prompt tuning specifically for heterophilic graphs.
Appendix E Experiment Results
In this section, we present the complete experimental results.
E.1 Parameter Sensitivity
@@
(i) Regarding the threshold for cross edges, we set the inner edge threshold at 0.2 (in line with the previous work Sun et al. [2023]) and vary the cross edge threshold from 0 to 1. The model is evaluated on three datasets with varying degrees of homophily: Cora, Wisconsin, and Empire. The results reveal a distinct homophily-driven heuristic in the selection of the optimal threshold: the optimal cross edge threshold for homophilic graphs (e.g., Pubmed) is higher than that for heterophilic graphs (e.g., Wisconsin and Empire). This systematic variation stems from our dual objectives of maintaining structural integrity and ensuring the effectiveness of the prompts. In homophilic graphs, where neighboring nodes tend to be similar, we opt for a higher $\tau_{cross}$ value to ensure that connected nodes have more similar features and preserve the graph’s homophilic nature. Conversely, in heterophilic graphs with more dissimilar neighboring nodes, we set a lower $\tau_{cross}$ to facilitate more diverse connections between prompt and original graph nodes.
@@ E.2 Runtime Analysis
E.3 Additional Accuracy Results
@@
GCN 0.6532 $\pm$0.0319 0.6506 $\pm$0.0185 0.5546 $\pm$0.0367 0.5405 $\pm$0.0358 0.4442 $\pm$0.0303 0.4187 $\pm$0.0220 0.2384 $\pm$0.0150 0.1911 $\pm$0.0634 0.1613 $\pm$0.0145 0.1491 $\pm$0.0707
GAT 0.6018 $\pm$0.0238 0.6091 $\pm$0.0201 0.5472 $\pm$0.0641 0.5381 $\pm$0.0595 0.4446 $\pm$0.0350 0.4162 $\pm$0.0332 0.2002 $\pm$0.0248 0.1827 $\pm$0.0371 0.1419 $\pm$0.0221 0.1367 $\pm$0.0208
BernNet 0.6494 $\pm$0.0280 0.6356 $\pm$0.0296 0.6584 $\pm$0.0216 0.6493 $\pm$0.0215 0.4888 $\pm$0.0197 0.4712 $\pm$0.0192 0.2581 $\pm$0.0074 0.2030 $\pm$0.0063 0.1215 $\pm$0.0280 0.1071 $\pm$0.0253
ChebNetII 0.6798 $\pm$0.0170 0.6752 $\pm$0.0202 0.6616 $\pm$0.0431 0.6554 $\pm$0.0410 0.4526 $\pm$0.0353 0.4304 $\pm$0.0328 0.2382 $\pm$0.0148 0.1970 $\pm$0.0047 0.0661 $\pm$0.0088 0.0551 $\pm$0.0054
TFE-GNN 0.3376 $\pm$0.0198 0.3286 $\pm$0.0177 0.4694 $\pm$0.0170 0.4590 $\pm$0.0155 0.2932 $\pm$0.0237 0.2874 $\pm$0.0203 0.2301 $\pm$0.0099 0.1939 $\pm$0.0014 0.3502 $\pm$0.0151 0.3155 $\pm$0.0182
DGI 0.3658 $\pm$0.1487 0.3187 $\pm$0.1251 0.4870 $\pm$0.0834 0.3743 $\pm$0.1073 0.2922 $\pm$0.0347 0.2235 $\pm$0.0405 0.1818 $\pm$0.0418 0.1485 $\pm$0.0237 0.1209 $\pm$0.0359 0.0710 $\pm$0.0188
GraphCL 0.5958 $\pm$0.0395 0.5603 $\pm$0.0486 0.5694 $\pm$0.0327 0.5576 $\pm$0.0396 0.4154 $\pm$0.0263 0.3741 $\pm$0.0356 0.2326 $\pm$0.0358 0.1853 $\pm$0.0158 0.1184 $\pm$0.0245 0.0961 $\pm$0.0150
SimGRACE 0.4400 $\pm$0.0321 0.4283 $\pm$0.0341 0.4416 $\pm$0.0424 0.4316 $\pm$0.0482 0.3664 $\pm$0.0392 0.3412 $\pm$0.0405 0.2064 $\pm$0.0239 0.1761 $\pm$0.0153 0.0878 $\pm$0.0097 0.0635 $\pm$0.0065
PolyGCL 0.6948 $\pm$0.0284 0.6655 $\pm$0.0221 0.6792 $\pm$0.0364 0.6782 $\pm$0.0357 0.5266 $\pm$0.0286 0.5010 $\pm$0.0224 0.2048 $\pm$0.0155 0.1874 $\pm$0.0081 0.0634 $\pm$0.0060 0.0528 $\pm$0.0029
GPPT 0.7026 $\pm$0.0172 0.6611 $\pm$0.0313 0.6540 $\pm$0.0452 0.6459 $\pm$0.0447 0.3854 $\pm$0.0617 0.3731 $\pm$0.0608 0.2345 $\pm$0.0156 0.1944 $\pm$0.0128 0.0750 $\pm$0.0037 0.0658 $\pm$0.0036
GPrompt 0.6578 $\pm$0.0323 0.6366 $\pm$0.0252 0.6230 $\pm$0.0428 0.6113 $\pm$0.0428 0.4568 $\pm$0.0269 0.4316 $\pm$0.0315 0.2269 $\pm$0.0252 0.1946 $\pm$0.0125 0.0652 $\pm$0.0071 0.0539 $\pm$0.0040
GPF-plus 0.6250 $\pm$0.0279 0.6160 $\pm$0.0251 0.6706 $\pm$0.0242 0.6534 $\pm$0.0250 0.4564 $\pm$0.0308 0.4441 $\pm$0.0322 0.2092 $\pm$0.0273 0.1826 $\pm$0.0177 0.0655 $\pm$0.0080 0.0476 $\pm$0.0055
All-in-One 0.5442 $\pm$0.0731 0.4308 $\pm$0.0410 0.6724 $\pm$0.0335 0.6600 $\pm$0.0330 0.4616 $\pm$0.0095 0.3638 $\pm$0.0158 0.2679 $\pm$0.0371 0.1819 $\pm$0.0263 0.0566 $\pm$0.0037 0.0472 $\pm$0.0028
ProNoG 0.5838 $\pm$0.0121 0.5564 $\pm$0.0203 0.5348 $\pm$0.0748 0.5242 $\pm$0.0855 0.2596 $\pm$0.0160 0.2466 $\pm$0.0172 0.2513 $\pm$0.0426 0.1963 $\pm$0.0102 0.1025 $\pm$0.0142 0.0784 $\pm$0.0090
HS-GPPT 0.7156 $\pm$0.0132 0.6915 $\pm$0.0166 0.6912 $\pm$0.0215 0.6910 $\pm$0.0199 0.5188 $\pm$0.0351 0.5043 $\pm$0.0223 0.2407 $\pm$0.0239 0.1972 $\pm$0.0123 0.3836 $\pm$0.0322 0.3520 $\pm$0.0194